The measure of the vertex angle of an isosceles triangle is 126. Find the measure of each base angle???
Answer by HyperBrain(694)

If the vertex angle is 126 degrees, then the sum of the two base angles is 180-126=54 degrees. Furthermore, the two base angles measures the same. Thus,

two angles =54
one angle =54/2=27 degrees
Therefore, the measure of one angle is 27 degrees!

Power up,
HyperBrain!